# SEW: A full-spectrum linear fitting with stellar population synthesis method Based on “Equivalent Widths spectrum”

### v1

@@ 3 Test with mock spectrum
Bias type Method S/N
5 15 30
No bias SEW (-0.12,0.27),(0.14,0.36) (-0.05,0.15),(0.06,0.22) (-0.03,0.1),(0.03,0.14)
pPXF+accurate (-0.04,0.15),(0.08,0.27) (-0.02,0.09),(0.04,0.16) (-0.01,0.06),(0.01,0.1)
pPXF+misfit (0.64,0.2),(0.01,0.34) (0.65,0.18),(-0.03,0.34) (0.65,0.18),(-0.04,0.35)
calibration bias SEW (-0.09,0.27),(0.13,0.36) (-0.05,0.14),(0.06,0.2) (-0.03,0.1),(0.04,0.15)
pPXF+accurate (-0.05,0.15),(0.11,0.27) (-0.02,0.09),(0.04,0.15) (-0.01,0.06),(0.02,0.11)
pPXF+curve (-0.4,0.19),(0.48,0.27) (-0.46,0.19),(0.5,0.22) (-0.47,0.19),(0.5,0.21)
pPXF+poly (-0.08,0.25),(0.65,0.29) (-0.07,0.18),(0.72,0.22) (-0.07,0.16),(0.73,0.21)
stitching bias SEW (-0.11,0.27),(0.13,0.35) (-0.05,0.14),(0.05,0.21) (-0.03,0.1),(0.02,0.14)
pPXF+accurate (-0.05,0.15),(0.1,0.27) (-0.02,0.09),(0.04,0.16) (-0.01,0.06),(0.02,0.1)
pPXF+curve (0.27,0.16),(-0.33,0.27) (0.31,0.13),(-0.43,0.22) (0.31,0.12),(-0.44,0.21)
pPXF+poly (-0.22,0.33),(0.12,0.37) (-0.12,0.18),(0.01,0.26) (-0.1,0.14),(-0.01,0.24)
Table 2: The statistical results for different bias and different fitting methods at various S/N levels in Figure 4. The results are shown as following forms: $(\overline{\Delta\log age_{L}},\sigma_{\Delta\log age_{L}}),(\overline{\Delta[Z/Z_{\odot}]},\sigma_{\Delta[Z/Z_{\odot}]})$.
3.4 statistical tests
[…]
Figure 4 displays the distributions of fitting deviations from true values $(\Delta\log age_{L},\Delta[Z/Z_{\odot}])$, with each subplot corresponding to specific test conditions annotated in the top-left corner (for direct comparisons between the true values and fitted values, see Appendix B). Different methodologies are colour-coded in the panel legend, with reference zero-points marked by ’X’ symbols. Table 2 quantifies these systematic offsets and their associated dispersions $(\overline{\Delta\log age_{L}},\sigma_{\Delta\log age_{L}}),(\overline{\Delta[Z/Z_{\odot}]},\sigma_{\Delta[Z/Z_{\odot}]})$ across various bias types and S/N levels.
Our analysis reveals comparable performance between the SEW method and pPXF with properly calibrated attenuation curves, both showing minor deviations within statistical uncertainty ranges (Table 2) in both unbiased calibration and biased calibration scenarios. The observed discrepancies likely originate from noise-induced spectral misidentification.
@@
Conversely, spectra exhibiting calibration systematics demonstrate substantial fitting deviations even when employing appropriate attenuation curves, provided that calibration offsets remain unaccounted for (red contours in lower panels of Figures 4 and B1). While standard practice addresses such systematics through high-order polynomial corrections, our testing demonstrates that even a 10th-degree polynomial fails to fully encapsulate calibration biases (orange contours in corresponding panels), resulting in residual deviations exceeding acceptable thresholds.
The fitting results without any prior calibration bias or unsuitable attenuation curve exhibit significant deviations from the intrinsic spectrum. We emphasize that the above deviations are specific to the predefined mock scenarios in this paper. In real observations, it is difficult to precisely assess various sources of bias. The systematic biases in stellar population synthesis results caused by such flux biases and attenuation curve choice bias are also challenging to estimate.
@@ 4 Discussion
For example, to suppress fluctuations of $\Delta\lambda=20$Å at $\lambda_{0}=5000$Å, the required penalty scale factor is about 2.5 ×$\times 10^{-10}$. This relationship highlights how the penalty scale factor $\sigma$ controls the trade-off between resolving fine spectral features (small $\sigma$) and smoothing out noise or rapid variations (large $\sigma$).
In this section, we will systematically validate the choice of $\sigma$, aiming to find an optimal balance that preserves the integrity of the continuous spectrum while effectively removing noise and unwanted features. This will ensure that the absorption line characteristics are accurately represented in the processed data. To explore this, we test how varying $\sigma$ affects the distribution of deviations between the fitted and true values of age and metallicity for the 1000 mock spectra. By analyzing these distributions, we can identify optimal $\sigma$ values that minimise the discrepancies between the fitted parameters and their known true values, thereby enhancing the accuracy and reliability of our method under different conditions. In this test, the $\sigma$ takes the values of $10^{-4}$, $10^{-6}$, $10^{-8}$, $10^{-10}$, $10^{-12}$ (corresponding to the characteristic scale $\Delta\lambda$ of about 500Å, 160Å, 50Å, 16Å, 5Å at $\lambda_{0}=5000$Å, respectively). Additionally, we apply an power-law attenuation curve with $E(B-V)$ = 0.3 and $\beta$ = 1.3.
The test results is shown in Figure 5. In this figure, we find that the performance of $\sigma$ in the range of $10^{-6}$ to $10^{-10}$ (corresponding to the characteristic scale $\Delta\lambda$ from about 160Å, to 16Å at $\lambda_{0}=5000$Å) is not significantly different. As shown from the second to the fourth column in Figure 5, the mean of $(\Delta\log age_{L},\Delta Z)$ does not vary substantially within this range. Slight deviations may be an amplification of intrinsic deviations in pPXF. As shown in the figure, the results of our method have the same trend but slightly larger deviations compared to pPXF. However, when $\sigma$ is $10^{-4}$ ($\Delta\lambda\sim 500$Å at $\lambda_{0}=5000$Å) and $10^{-12}$ ($\Delta\lambda\sim 5$Å at $\lambda_{0}=5000$Å), the results of our method exhibit a much larger and off-trend deviation.
@@ Appendix A Attenuation Curve
Figure A1: This figure illustrates the attenuation curves obtained using the SEW method under different reddening ( $E(B-V)$ ) and signal-to-noise ratio (S/N) conditions. Each row corresponds to a specific $E(B-V)$ value (0.0, 0.1, 0.3, and 0.5), and the left three columns show the attenuation curves for different S/N ratios (5, 15, and 30). In each panel of the left three columns, the density plot represents the distribution of 1000 mock fits, with the blue dashed line indicating the mean value and the black solid line representing the true value. The rightmost column displays the parameter distributions obtained by fitting the attenuation curves in the corresponding rows using a power law. Different colours in the rightmost column represent different S/N ratios: red for S/N = 5, magenta for S/N = 15, and blue for S/N = 30. The circle dot represents the initial $E(B-V)$ and $\beta$ and different colour “X" mark represent the mean $E(B-V)$ and $\beta$ at the corresponding S/N.
This appendix validates attenuation curve recovery of the SEW method through controlled simulations using synthetic galaxy spectra. We systematically vary dust reddening $E(B-V)\in[0,0.1,0.3,0.5]$ (fixed $\beta=1.3$) and observational noise levels ($\mathrm{S/N}=5,15,30$), intentionally excluding calibration bias for direct benchmarking. Figure A1 demonstrates consistent recovery of attenuation curves across all tested $E(B-V)$ values, with mean reconstructed curves (blue dashed lines) closely matching true values (black lines) despite increasing dispersion at lower S/N. The rightmost KDE plots reveal tight clustering around true parameters at S/N $\geq$15, while S/N=5 shows broader distributions. Notably, $E(B-V)$ dispersion remains constant across S/N levels due to combined uncertainties from stellar population fitting and random noise, whereas $\beta$ dispersion depends on $E(B-V)$ magnitude - weaker constraints emerge for lower $E(B-V)$ where attenuation curves flatten. These results confirm the robustness across diverse observational conditions, though precision decreases significantly at S/N=5.
@@
Figure B1: The age (top) and metallicity (bottom) comparison between fitted and True means for different methods and various S/N and flux bias settings. Each subplot displays a specific test condition, labeled in the upper left corner of the plot. Different methodologies are colour-coded in the panel legend.

### v2

@@ 3.3 Flux bias robustness of SEW
Bias type Method S/N=5 S/N=15 S/N=30
$\Delta\log\text{age}_{L}$ $\Delta[Z/Z_{\odot}]$ $\Delta\log\text{age}_{L}$ $\Delta[Z/Z_{\odot}]$ $\Delta\log\text{age}_{L}$ $\Delta[Z/Z_{\odot}]$
No bias SEW $0.12\pm 0.27$ $14\pm 0.36 0.0$$5\pm 0.15 0.0$$6\pm 0.22 0.0$$3\pm 0.1 0.0$$03\pm 0.14$
pPXF+accurate $04\pm 0.15 0.0$$8\pm 0.27 0.0$$2\pm 0.09 0.0$$4\pm 0.16 0.0$$1\pm 0.06 0.0$$01\pm 0.1$
pPXF+misfit $64\pm 0.2 0.0$$01\pm 0.34$ $65\pm 0.18 0.0$$03\pm 0.34$ $65\pm 0.18 0.0$$04\pm 0.35$
Calibration bias SEW $0.09\pm 0.27$ $13\pm 0.36 0.0$$5\pm 0.14 0.0$$6\pm 0.2 0.0$$3\pm 0.1 0.0$$04\pm 0.15$
pPXF+accurate $0.05\pm 0.15$ $11\pm 0.27 0.0$$2\pm 0.09 0.0$$4\pm 0.15 0.0$$1\pm 0.06 0.0$$02\pm 0.11$
pPXF+curve $0.4\pm 0.19$ $0.48\pm 0.27$ $0.46\pm 0.19$ $0.5\pm 0.22$ $0.47\pm 0.19$ $0.5\pm 0.21$
pPXF+poly $0.35\pm 0.24$ $0.66\pm 0.28$ $0.41\pm 0.18$ $0.71\pm 0.23$ $0.42\pm 0.17$ $0.74\pm 0.22$
Stitching bias SEW $0.11\pm 0.27$ $13\pm 0.35 0.0$$5\pm 0.14 0.0$$5\pm 0.21 0.0$$3\pm 0.1 0.0$$02\pm 0.14$
pPXF+accurate $0.05\pm 0.15$ $1\pm 0.27 0.0$$2\pm 0.09 0.0$$4\pm 0.16 0.0$$1\pm 0.06 0.0$$02\pm 0.1$
pPXF+curve $0.27\pm 0.16$ $0.33\pm 0.27$ $0.31\pm 0.13$ $0.43\pm 0.22$ $0.31\pm 0.12$ $0.44\pm 0.21$
pPXF+poly $07\pm 0.25 0.0$$3\pm 0.33 0.0$$01\pm 0.13$ $19\pm 0.22 0.0$$03\pm 0.10$ $0.23\pm 0.17$
Table 2: The statistical results for different bias types and fitting methods at various S/N levels (corresponding to Figure 4). The results are formatted as $\overline{\Delta\log\text{age}_{L}}\pm\sigma_{\Delta\log\text{age}_{L}}$ and $\overline{\Delta[Z/Z_{\odot}]}\pm\sigma_{\Delta[Z/Z_{\odot}]}$.
3.4 Statistical Results
To quantitatively evaluate the SEW method, we synthesize 1,000 mock spectra following the procedures detailed in Section 3.1. We conduct systematic evaluations of the method reliability across varying signal-to-noise ratios (S/N = 5, 15, 30) under three distinct flux calibration scenarios: unbiased flux calibration, calibration-induced bias, and spectral stitching artefacts (as defined in Sections 3.2 and 3.3). The attenuation model employs a power law parameterized withE(BV)E(B-V)= 0.3 and $\beta=1.3$. For each combination of flux calibration condition and S/N level, we apply identical fitting methodologies to those described in the aforementioned sections. In biased conditions (calibration bias and stitching artefacts), we additionally implement a pPXF-based approach incorporating “accurate curve with systematic error correction” to assess potential improvement.
Figure 4 displays the distributions of fitting deviations from true values $(\Delta\log\text{age}_{L},\Delta[Z/Z_{\odot}])$, with each subplot corresponding to specific test conditions annotated in the top-left corner (for direct comparisons between the true values and fitted values, see Appendix B). Different methodologies are colour-coded in the panel legend, with reference zero-points marked by ’X’ symbols. Table 2 quantifies these systematic offsets and their associated dispersions $(\overline{\Delta\log\text{age}_{L}},\sigma_{\Delta\log\text{age}_{L}}),(\overline{\Delta[Z/Z_{\odot}]},\sigma_{\Delta[Z/Z_{\odot}]})$ across various bias types and S/N levels.
Our analysis reveals comparable performance between the SEW method and pPXF with properly calibrated attenuation curves, both showing minor deviations within statistical uncertainty ranges (Table 2) in both unbiased calibration and biased calibration scenarios. The observed discrepancies likely originate from noise-induced spectral misidentification. In particular, misidentification of error signals as metal absorption lines leads to overestimation of metallicity at low metallicity with low S/N (bottom panel of Figure B1)
@@ 3.4 Statistical Results
Conversely, spectra exhibiting calibration systematics demonstrate substantial fitting deviations even when employing appropriate attenuation curves, provided that calibration offsets remain unaccounted for (red contours in lower panels of Figures 4 and B1). Of course, higher calibration precision naturally yields smaller residual biases, eventually degrading to a bias-free case. While standard practice addresses such systematics through high-order polynomial corrections, our testing demonstrates that even a 10th-degree polynomial fails to fully encapsulate calibration biases (orange contours in corresponding panels) in our cases, resulting in residual deviations exceeding acceptable thresholds. This is because polynomial corrections need to take into account both the calibration bias and the existence of extinction. This limitation stems from intrinsic challenges in multiplicative polynomial calibration: the arbitrary selection of polynomial order introduces artificial effects, where insufficient polynomial degrees fail to resolve intermediate-scale spectral deviations, while excessive degrees induce non-linear parameter degeneracies (obscuring physical parameters amidst numerous polynomial coefficients) and escalate computational demands through prolonged optimisation processes.
The fitting results without any prior calibration bias correction or appropriate attenuation curve selection exhibit significant deviations from the intrinsic spectrum. We emphasize that these deviations are specific to the predefined mock scenarios in this paper. While these deviations are specific to our predefined mock scenarios, we emphasize that the adopted about $5\%$ flux bias serves as a deliberately strong assumption to amplify the potential impact of systematic errors for illustrative purposes. In real observations, it is difficult to precisely assess various sources of bias. The systematic biases in stellar population synthesis results caused by such flux biases and attenuation curve choice bias are also challenging to estimate. Although our simplified bias model highlights first-order effects, we caution that real systematic uncertainties likely arise from compounded effects, including degeneracies between stellar population parameters, attenuation curve flexibility, and wavelength-dependent flux calibration errors, whose collective behaviour requires more sophisticated error propagation frameworks. For instance, recently detected intermediate-scale structures in extinction curves (FWHM $\sim$ 100 Å; Massa et al. 2020; Zhang et al. 2024; Green et al. 2024) highlights how insufficient polynomial order compromises measurement accuracy of spectral features.
@@ 4 Discussion
For example, to suppress fluctuations of $\Delta\lambda=20$Å at $\lambda_{0}=5000$Å, the required penalty scale factor is about 2.5 $\times 10^{-10}$. This relationship highlights how the penalty scale factor $\sigma$ controls the trade-off between resolving fine spectral features (small $\sigma$) and smoothing out noise or rapid variations (large $\sigma$).
In this section, we will systematically validate the choice of $\sigma$, aiming to find an optimal balance that preserves the integrity of the continuous spectrum while effectively removing noise and unwanted features. This will ensure that the absorption line characteristics are accurately represented in the processed data. To explore this, we test how varying $\sigma$ affects the distribution of deviations between the fitted and true values of age and metallicity for the 1000 mock spectra. By analyzing these distributions, we can identify optimal $\sigma$ values that minimise the discrepancies between the fitted parameters and their known true values, thereby enhancing the accuracy and reliability of our method under different conditions. In this test, the $\sigma$ takes the values of $10^{-4}$, $10^{-6}$, $^{-8}, 1010$$^{-10}$, $10^{-12}$ (corresponding to the characteristic scale $\Delta\lambda$ of about 500Å, 160Å, 50Å, 16Å, 5Å at $\lambda_{0}=5000$Å, respectively). Additionally, we apply an power-law attenuation curve with E(BV)E(B-V) = 0.3 and $\beta$ = 1.3.
The test results is shown in Figure 5. In this figure, we find that the performance of $\sigma$ in the range of $10^{-6}$ to $10^{-10}$ (corresponding to the characteristic scale $\Delta\lambda$ from about 160Å, to 16Å at $\lambda_{0}=5000$Å) is not significantly different. As shown from the second to the fourth column in Figure 5, the mean of $(\Delta\log\text{age}_{L},\Delta[Z/Z_{\odot}])$ does not vary substantially within this range. Slight deviations may be an amplification of intrinsic deviations in pPXF. As shown in the figure, the results of our method have the same trend but slightly larger deviations compared to pPXF. However, when $\sigma$ is $10^{-4}$ ($\Delta\lambda\sim 500$Å at $\lambda_{0}=5000$Å) and $10^{-12}$ ($\Delta\lambda\sim 5$Å at $\lambda_{0}=5000$Å), the results of our method exhibit a much larger and off-trend deviation.
@@ Appendix A Attenuation Curve
Figure A1: This figure illustrates the attenuation curves obtained using the SEW method under different reddening ( E(BV)E(B-V) ) and signal-to-noise ratio (S/N) conditions. Each row corresponds to a specific E(BV)E(B-V) value (0.0, 0.1, 0.3, and 0.5), and the left three columns show the attenuation curves for different S/N ratios (5, 15, and 30). In each panel of the left three columns, the density plot represents the distribution of 1000 mock fits, with the blue dashed line indicating the mean value and the black solid line representing the true value. The rightmost column displays the parameter distributions obtained by fitting the attenuation curves in the corresponding rows using a power law. Different colours in the rightmost column represent different S/N ratios: red for S/N = 5, magenta for S/N = 15, and blue for S/N = 30. The circle dot represents the intrinsic E(BV)E(B-V) and $\beta$ and different colour “X" mark represent the mean E(BV)E(B-V) and $\beta$ at the corresponding S/N.
This appendix validates attenuation curve recovery of the SEW method through controlled simulations using synthetic galaxy spectra. We systematically vary dust reddening $E(B-V)\in[0,0.1,0.3,0.5]$ (fixed $\beta=1.3$) and observational noise levels ($\mathrm{S/N}=5,15,30$), intentionally excluding calibration bias for direct benchmarking. Figure A1 demonstrates consistent recovery of attenuation curves across all tested E(BV)E(B-V) values, with mean reconstructed curves (blue dashed lines) closely matching true values (black lines) despite increasing dispersion at lower S/N. The rightmost KDE plots reveal tight clustering around true parameters at S/N $\geq$15, while S/N=5 shows broader distributions. Notably, E(BV)E(B-V) dispersion remains constant across S/N levels due to combined uncertainties from stellar population fitting and random noise, whereas $\beta$ dispersion depends on E(BV)E(B-V) magnitude - weaker constraints emerge for lower E(BV)E(B-V) where attenuation curves flatten. These results confirm the robustness across diverse observational conditions, though precision decreases significantly at S/N=5.
@@
Figure B1: The age (top) and metallicity (bottom) comparison between fitted and True means for different methods and various S/N and flux bias settings. Each subplot displays a specific test condition, labeled in the upper left corner of the plot. Different methodologies are colour-coded in the panel legend, with matching colored error bars indicating binned average measurements along the parameter space.